OpenStudy (anonymous):

solve for t .... I got 10 3t – 6t + 8 = 7t – 2 A. 1 B. 4 C. 10 D. 8

OpenStudy (anonymous):

You cant find t if i does not tell you itself

OpenStudy (yttrium):

No. It's not 10 because it is 10t = 10. Hence, what is your t??

OpenStudy (anonymous):

3t – 6t + 8 = 7t – 2 -3t+8=7t-2 8=10t-2 10=10t ?=t

OpenStudy (mony01):

A. 1

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thanks i got A too